OpenStudy (anonymous):

Solve the linear system of equations. 7x-4y=2 y=2x+3

OpenStudy (pinksapphire):

First put 2x+3 into the first equation like this: 7x-4(2x+3)=2

OpenStudy (anonymous):

7x-4(2x+3)=2 7x-8x+12=2 ??? -1x-12=2 -1x=14 did i do that right?

OpenStudy (pinksapphire):

Then you distribute: 7x-8x-12=2

OpenStudy (anonymous):

x=-14

OpenStudy (pinksapphire):

Check again, -12 is negative because -4*3 turns into a negative 12

OpenStudy (pinksapphire):

Correct!

OpenStudy (pinksapphire):

Plug in -14 into y=2x+3

OpenStudy (pinksapphire):

y=2(-14)+3

OpenStudy (anonymous):

y=2(-14)+3 y=-28+3=-25

OpenStudy (pinksapphire):

Yes.

OpenStudy (anonymous):

Ok thanks